### v1

# AVSegFormer: Audio-Visual Segmentation with Transformer

Abstract.
The combination of audio and vision has long been a topic of interest in the multi-modal community. Recently, a new audio-visual segmentation (AVS) task has been introduced, aiming to locate and segment the sounding objects in a given video. This task demands audio-driven pixel-level scene understanding for the first time, posing significant challenges. In this paper, we propose AVSegFormer, a novel framework for AVS tasks that leverages the transformer architecture. Specifically, we introduce audio queries and learnable queries into the transformer decoder, enabling the network to selectively attend to interested visual features. Besides, we present an audio-visual mixer, which can dynamically adjust visual features by amplifying relevant and suppressing irrelevant spatial channels. Additionally, we devise an intermediate mask loss to enhance the supervision of the decoder, encouraging the network to produce more accurate intermediate predictions. Extensive experiments demonstrate that AVSegFormer achieves state-of-the-art results on the AVS benchmark. The code is available at https://github.com/vvvb-github/AVSegFormer.
1. Introduction
Audio and vision are closely intertwined modalities that play crucial roles in the perception of the world. For example, we rely heavily on auditory and visual cues to comprehend and navigate our surroundings. These underlying connections motivated many audio-visual tasks, such as audio-visual correspondence [4, 5], audio-visual event localization [29, 31, 46], audio-visual video parsing [45, 51, 30], and sound source localization [4, 5, 43]. However, due to the lack of pixel-level annotations for these tasks, they are often limited to the frame/temporal level, which makes them become audio-informed image classification problems eventually.
Recently, a novel audio-visual segmentation (AVS) [56] task has been proposed, which aims to segment sounding objects from video frames corresponding to a given audio. It includes three sub-tasks: single sound source segmentation (S4), multiple sound source segmentation (MS3), and audio-visual semantic segmentation (AVSS). Figure 1 illustrates the objectives of the three sub-tasks. Compared to previous audio-visual tasks, the AVS task presents a greater challenge as it requires the network to not only locate the audible frames but also delineate the shape of the sounding objects [57, 56]. This demands the alignment of multiple modalities and necessitates a detailed understanding of the scenarios. These unique characteristics of the AVS task make many existing methods [7, 39, 36] designed for other audio-visual tasks may be sub-optimal on AVS. Therefore, designing new methods tailored for AVS becomes necessary.
For the brand-new AVS task, AVSBench [56] designed a strong baseline method that achieves state-of-the-art audio-visual segmentation performance. Figure 2(a) illustrates its network architecture, which incorporates a modality fusion module before the convolution-based decoder (i.e., Semantic FPN [26]) to enable audio-visual segmentation. This design is simple yet effective, but some inherent flaws of convolution still limit it. (1) The effective receptive field of convolutions is relatively small. Even with a deep decoder, the audio feature still cannot capture long-range visual dependencies, which restricts its performance. (2) Convolution is an operator with static weights, which is difficult to provide different visual features conditioned by the input audio.
To remedy these issues, we propose AVSegFormer, a novel framework for audio-visual segmentation with transformers. The brief architecture is shown in Figure 2(b). First, we introduce audio queries along with learnable queries into the segmentation decoder, enabling the network to selectively attend to interested visual features. Second, we present an audio-visual mixer, which can dynamically adjust visual features by amplifying relevant and suppressing irrelevant spatial channels, allowing visual features to adapt to diverse audio features. Third, an intermediate mask loss is designed to enhance the supervision of the decoder, which encourages the network to produce more accurate intermediate predictions and helps refine the final segmentation outputs.
We evaluate AVSegFormer on the three sub-tasks of AVS, including S4, MS3, and AVSS, with widely-used backbones ResNet-50 [20] and PVTv2 [49]. Our experimental results demonstrate that AVSegFormer significantly outperforms existing state-of-the-art methods, such as LGVT [55], SST [17], iGAN [37], and AVSBench baseline [57]. Specifically, AVSegFormer-R50 achieves 76.45, 49.53, and 24.93 mIoU on S4, MS3, and AVSS, surpassing the AVSBench-R50 by 3.66, 1.65, and 4.75 mIoU, respectively. Furthermore, using PVTv2 as the backbone, AVSegFormer yields consistently higher segmentation performance on all three sub-tasks, setting new state-of-the-art records of 82.06, 58.36, and 36.66 mIoU.
Overall, our contributions to this work are four-fold.
We employ audio queries and learnable queries to selectively attend to relevant visual features, overcoming the limitations of previous convolutional-based approaches.
We design an audio-visual mixer, which can amplify relevant visual features and suppress irrelevant visual features in response to audio cues.
We present an intermediate mask loss that provides additional supervision, encouraging more accurate intermediate results and improving final prediction.
We conduct extensive experiments on three sub-tasks of AVS. These results demonstrate that our method achieves state-of-the-art performance on the AVS benchmark.
2. Related Works
2.1. Multi-Modal Tasks
In recent years, multi-modal tasks have gained significant attention in the research community due to their potential for advancing the understanding of complex real-world scenarios. These tasks aim to exploit the complementary information from different modalities, such as vision, audio, and text, to improve the overall performance of various applications. Among these, text-visual tasks have attracted considerable interest from researchers. Numerous works focus on related tasks, such as visual question answering [3, 50, 2, 9] and visual grounding [13, 15, 25, 32]. In addition to text-visual tasks, audio-visual tasks are emerging as hot spots in the multi-modal field. Several related tasks include audio-visual correspondence [4, 5], audio-visual event localization [29, 31, 46], and sound source localization [4, 5, 43]. Concurrently, some works [59, 48] have proposed unified architectures capable of handling nearly all types of modalities using a consistent format for embeddings.
Most of these works are based on transformer architecture [47], which demonstrates a strong cross-modal capability. Their achievements highlight the reliability of transformers in the multi-modal field. As a recently proposed multi-modal task, audio-visual segmentation (AVS) [57, 56] shares many commonalities with the aforementioned tasks. The pioneering works in these areas have significantly inspired our research of AVSegFormer.
2.2. Sound Source Localization
Sound source localization (SSL) is an important problem in the audio-visual multi-modal community and is also the most related task to audio-visual segmentation. It aims to locate the regions for sound sources in a video sequence, but the results are usually represented as heat maps. The major challenge in the SSL problem is dealing with multiple sound sources. In prior arts, Hu et al. [22] divided audio-visual features into multiple clustering centers and used center distance as a supervised signal to locate paired audio-visual information. Qian et al. [39] trained an audio-visual correspondence model to extract coarse feature representations of audio and visual signals and used Grad-CAM [42] to locate specific categories of features. Hu et al. [23] adopted a two-stage approach by first learning audio-visual semantics under a single sound source and then using this knowledge to help locate multiple sound sources. Besides, Rouditchenko et al. [41] tackled the problem by unraveling the concept of categories in neural networks.
Although these SSL methods indicate which areas in the image emit sound, they do not clearly depict the shape of the object, which is another challenge in the AVS task. Furthermore, the above methods all rely on unsupervised learning when capturing the shape of the detected object, which may result in inaccurate localization. Nevertheless, handling multiple sound sources in SSL methods offers valuable insights and references for our work.
2.3. Vision Transformer
During the past few years, Transformer [47] has experienced rapid development in natural language processing. Following this success, the Vision Transformer (ViT) [16] emerged, bringing the transformer into the realm of computer vision and yielding impressive results. Numerous works [33, 49, 10] have built upon ViT, leading to the maturation of vision transformers. As the performance of vision transformers continues to advance, they are increasingly replacing CNNs as the mainstream paradigm in the field of computer vision, especially in object detection and image segmentation tasks.
Carion et al. [6] proposed the DETR model and designed a novel bipartite matching loss based on the transformer architecture, paving the way for new research directions in vision transformers. Subsequently, improved frameworks such as Deformable DETR [58] and DINO [54] are proposed, introducing mechanisms like deformable attention and denoise training. These arts take vision transformers to new heights. The remarkable performance of vision transformers has also inspired us to apply this paradigm to AVS tasks, anticipating further advancements in the field.
2.4. Image Segmentation
Image segmentation is a critical visual task that involves partitioning an image into distinct segments or regions. It includes three different tasks: instance segmentation, semantic segmentation, and panoptic segmentation. Instance segmentation predicts the mask of each object instance and its corresponding category, while semantic segmentation needs to classify each pixel in the image into different semantic categories. Panoptic segmentation unifies instance and semantic segmentation tasks and predicts the mask of each object instance or background segment. Early research proposed specialized models for these tasks, such as Mask R-CNN [19] and HTC [8] for instance segmentation, or FCN [34] and U-Net [40] for semantic segmentation. After panoptic segmentation was proposed, some related research [26, 52, 28] were conducted and designed universal models for both tasks.
The recent introduction of the transformer has led to the development of new models that can unify all the segmentation tasks. Mask2Former [11] is one such model that introduces mask attention into the transformer and improves MaskFormer [12]. Mask DINO [27] is a unified transformer-based framework for both detection and segmentation. Recently, OneFormer [24] presented a new multi-task universal image segmentation framework with transformers. These models have brought image segmentation to a new level, demonstrating the potential of transformer architecture in vision tasks. Considering that the AVS task involves segmentation, these methods have significantly contributed to our work.
3. METHOD
3.1. Overall Architecture
Figure 3 illustrates the overall architecture of our method. In contrast to previous CNN-based segmentation methods [57, 56], we design a query-based framework to leverage the transformer architecture. Specifically, our model combines audio queries with learnable queries, allowing it to adjust its focus on visual features dynamically. Additionally, we design an audio-visual mixer and an intermediate mask loss as auxiliary components. The audio-visual mixer aids in amplifying relevant features and suppressing irrelevant ones, while the intermediate mask loss helps supervise intermediate predictions for enhancing performance.
The pipeline of AVSegFormer consists of four stages. First, a visual backbone and an audio backbone are employed to extract features from video and audio frames, respectively. Second, the transformer encoder refines the visual features and generates an initial mask feature, which serves as the basis for predicting the final mask. Third, an audio-visual mixer is utilized to amplify feature channels relevant to sounding objects while suppressing those that are irrelevant. Lastly, the transformer decoder incorporates audio queries and learnable queries, capturing richer features about sounding objects and predicting the final mask.
3.2. Multi-Modal Representation
Visual encoder. We follow the feature extraction process adopted in previous methods [57, 56], which involves using a visual backbone and an audio backbone to extract video and audio features, respectively. For videos, the dataset provides pre-extracted frame images, making the process similar to image feature extraction. Specifically, the input video frames are denoted as , in which denotes the number of frames. Then, we use a visual backbone (e.g., ResNet-50 [20]) to extract hierarchical visual features , which can be written as:
in which and . represents the number of output channels of the stage of the visual backbone. In other words, is a list of multi-scale features, where each feature map is half resolution of the previous one.
Audio encoder. The process of audio feature extraction follows the VGGish [21] method. Initially, the audio is resampled to 16kHz mono audio , where is related to the audio duration. Then, we perform a short-time Fourier transform to obtain a mel spectrum. The mel spectrum is calculated by mapping the spectrum to a 64th-order mel filter bank and subsequently fed into the VGGish model to obtain the audio features , where represents the number of frames and defaults to 128 in the VGGish.
Feature transformation. For the convenience of subsequent usage, we use multiple linear layers to unify the number of channels for all features. Specifically, all features extracted by the visual backbone and audio backbone are transformed into dimensions, which also equals the embedding dimension of the transformer encoder and decoder. Typically, is set to 256 by default.
3.3. Transformer Encoder
The transformer encoder is responsible for multi-scale feature fusion and mask feature generation. Specifically, we collect the visual features of three resolutions (i.e., 1/8, 1/16, and 1/32), and then flatten and concatenate them as the input for the transformer encoder. After that, the output features of the transformer encoder are reshaped to their original shapes, and the 1/8-scale features are taken out separately and 2 upsampled. Finally, we add the upsampled features to the 1/4-scale features from the visual backbone and obtain the initial mask features , where , and is the embedding dimension of the transformer encoder and decoder.
3.4. Audio-Visual Mixer
As illustrated in Figure 3, the segmentation mask is generated based on the mask feature, which plays a crucial role in the final prediction results. However, since the audio features can vary widely, a static network may not be able to capture all of the relevant information. This limitation may hinder the model’s ability to identify potential sounding objects accurately.
To address this issue, we propose an audio-visual mixer as shown in Figure 4(b). The design of this module is based on channel attention, which allows the model to selectively amplify or suppress different visual channels depending on the audio input, improving its ability to capture complex audio-visual relationships. Specifically, the mixer learns a set of weights through audio-visual cross-attention, and applies them to highlight the relevant channels. The whole process can be represented as follows:
Here, and represent the input audio features and the initial mask features, and denotes the mixed mask features. Besides, means the number of attention heads, which is set to 8 by default following common practice.
3.5. Transformer Decoder
Audio query. The transformer decoder is designed to learn semantic-rich features of the sounding objects. We repeat the audio feature to the number of queries , and employ it as the audio queries . As the decoding process continues, the queries continuously aggregate visual information from the encoder’s outputs. The output queries ultimately combine the auditory and visual modalities and contain richer features of the sounding objects.
Learnable query. To enhance the model’s ability to capture in-depth information about the sounding object, we add learnable queries to the audio queries . The learnable queries enhance our model’s adaptability for various AVS tasks and datasets. Specifically, it enables the model to learn dataset-level contextual information, and adjust the attention allocated to different target categories. Furthermore, learnable queries empower the model with a more robust capability to process target semantics, ultimately improving segmentation accuracy.
Mask generation. To generate the segmentation masks, we multiply the mask feature obtained from the audio-visual mixer with the output queries from the decoder. Then, an MLP is used to integrate different channels. Additionally, we introduce a residual connection to ensure that the fusion of auditory information does not result in excessive loss of visual information. Finally, the model predicts the segmentation mask through a simple linear layer:
Here, represents the MLP process, and means the linear layer. The output is the predicted segmentation mask, with the dimension denotes the number of semantic classes.
3.6. Loss Function
Intermediate mask loss. With the introduction of the deep transformer decoder, achieving satisfying performance only by supervising the final prediction becomes more challenging. To address this issue, we design an intermediate mask loss as the auxiliary loss, which supervises each layer of the transformer decoder.
Specifically, the intermediate mask loss is based on the cross-attention operation of each decoder layer. First, we utilize the input queries and keys to compute the attention map . Then, we multiply the attention map with the output queries of the current decoder layer and feed it through a simple FPN [44] to predict a mask , as shown in Figure 5(c). Finally, we calculate the Dice loss [38] between this predicted mask and the ground truth to obtain the intermediate mask loss .
Total loss. The total loss function comprises two parts: IoU loss and mask loss. As shown in Figure 5(a), the IoU loss is calculated by comparing the final segmentation mask with the ground truth. Here, we also use Dice loss [38] for supervision. Considering that in AVS tasks, the proportion of segmented objects occupying the entire image is relatively small, the model can better focus on the foreground and reduce interference from the background by using Dice loss. Thus, the total loss of our method is as follows:
Here, represents the number of decoder layers, and is a coefficient that controls the effect of the auxiliary loss. We choose in our experiments as it performs best.
| Method | Backbone | S4 | MS3 | AVSS | Reference | |||
| F-score | mIoU | F-score | mIoU | F-score | mIoU | |||
| LVS [7] | ResNet-50 [20] | 51.0 | 37.94 | 33.0 | 29.45 | CVPR’2021 | ||
| MSSL [39] | ResNet-18 [20] | 66.3 | 44.89 | 36.3 | 26.13 | ECCV’2020 | ||
| 3DC [36] | ResNet-34 [20] | 75.9 | 57.10 | 50.3 | 36.92 | 21.6 | 17.27 | BMVC’2020 |
| SST [17] | ResNet-101 [20] | 80.1 | 66.29 | 57.2 | 42.57 | CVPR’2021 | ||
| AOT [53] | Swin-B [33] | 31.0 | 25.40 | NeurIPS’2021 | ||||
| iGAN [37] | Swin-T [33] | 77.8 | 61.59 | 54.4 | 42.89 | ArXiv’2022 | ||
| LGVT [55] | Swin-T [33] | 87.3 | 74.94 | 59.3 | 40.71 | NeurIPS’2021 | ||
| AVSBench-R50 [56] | ResNet-50 [20] | 84.8 | 72.79 | 57.8 | 47.88 | 25.2 | 20.18 | ECCV’2022 |
| AVSegFormer-R50 (ours) | ResNet-50 [20] | 85.9 | 76.45 | 62.8 | 49.53 | 29.3 | 24.93 | |
| AVSBench-PVTv2 [56] | PVTv2 [49] | 87.9 | 78.74 | 64.5 | 54.00 | 35.2 | 29.77 | ECCV’2022 |
| AVSegFormer-PVTv2 (ours) | PVTv2 [49] | 89.9 | 82.06 | 69.3 | 58.36 | 42.0 | 36.66 | |
| S4 | MS3 | |||
| mIoU | F | mIoU | F | |
| 1 | 81.1 | 88.6 | 57.1 | 67.7 |
| 100 | 81.4 | 88.9 | 57.5 | 68.2 |
| 200 | 81.9 | 89.6 | 58.0 | 68.4 |
| 300 | 82.0 | 89.9 | 58.3 | 69.3 |
| learnable | S4 | MS3 | ||
| queries | mIoU | F | mIoU | F |
| ✓ | 82.0 | 89.9 | 58.3 | 69.3 |
| 81.8 | 89.6 | 57.2 | 67.9 | |
| mixer | S4 | MS3 | ||
| mIoU | F | mIoU | F | |
| 81.4 | 89.0 | 57.4 | 67.9 | |
| CRA | 81.7 | 89.7 | 57.8 | 68.2 |
| CHA | 82.0 | 89.9 | 58.3 | 69.3 |
| mask | S4 | MS3 | ||
| reformer | mIoU | F | mIoU | F |
| 81.1 | 88.4 | 57.3 | 68.3 | |
| AR | 81.7 | 89.6 | 57.9 | 68.8 |
| QR | 82.0 | 89.9 | 58.3 | 69.3 |
4. Experiments
4.1. Dataset
AVSBench-Object [57] is an audio-visual dataset specifically designed for the audio-visual segmentation task, containing pixel-level annotations. The videos are downloaded from YouTube and cropped to 5 seconds, with one frame per second extracted for segmentation. The dataset includes two subsets: a semi-supervised single sound source subset for single sound source segmentation (S4), and a fully supervised multi-source subset for multiple sound source segmentation (MS3). S4 subset: The S4 subset contains a total of videos, with videos for training, for validation, and for testing. The target objects cover 23 different categories, including humans, animals, vehicles, and musical instruments. Besides, this subset is trained in a semi-supervised manner, where each video contains five frames, but only the first frame is annotated. MS3 subset: The MS3 subset includes 424 videos, with 286 training, 64 validation, and 64 testing videos, covering the same categories as the S4 subset. Additionally, unlike S4, the MS3 subset is full-supervised with all five frames annotated in training.
AVSBench-Semantic [56] is an extension of the AVSBench-Object dataset, which offers additional semantic labels that are not available in the original AVSBench-Object dataset. It is designed for audio-visual semantic segmentation (AVSS). In addition, the videos in AVSBench-Semantic are longer, with a duration of 10 seconds, and 10 frames are extracted from each video for prediction. It combines semi-supervised and fully-supervised manners, with labels provided for the first and first five frames of videos inherited from the S4 and MS3 subsets, respectively. For the newly collected 7,000 videos, a complete 10-frame label is provided. Overall, the AVSBench-Semantic dataset has increased in size by approximately three times compared to the original AVSBench-Object dataset, with 8,498 training, 1,304 validation, and 1,554 test videos.
Metric. These benchmark datasets employ mean intersection over union (mIoU) and F-score as the evaluation metrics.
4.2. Implementation Details
Our method is evaluated on three AVS sub-tasks, including single sound source segmentation (S4), multiple sound source segmentation (MS3), and audio-visual semantic segmentation (AVSS).
Details. We train our AVSegFormer models for the three sub-tasks using an NVIDIA V100 GPU. Consistent with previous works [57, 56], we employ AdamW [35] as the optimizer, with a batch size of 4 and an initial learning rate of . All video frames are resized to 224 224 resolution. For the S4 and MS3 subsets, each video contains 5 frames, while each video in AVSS contains 10 frames. Since the MS3 subset is quite small, we train it for 60 epochs, while the S4 and AVSS subsets are trained for 30 epochs. The encoder and decoder in our AVSegFormer both are comprised of 6 layers with an embedding dimension of 256. We set the coefficient of the proposed intermediate mask loss to 0.1 for the best performance. More detailed training settings can be found in Table 3.
4.3. Comparison with Prior Arts
To verify the effectiveness of our method, we conducted a comprehensive comparison between our AVSegFormer and existing methods on the AVS benchmark [57, 56]. For fairness, we employ the ImageNet-1K [14] pre-trained ResNet-50 [20] or PVTv2 [49] as the backbone to extract visual features, and the AudioSet [18] pre-trained VGGish [21] to extract audio features.
Comparison with methods from related tasks. Firstly, we compare our AVSegFormer with state-of-the-art methods from three AVS-related tasks, including sound source localization (LVS [7] and MSSL [39]), video object segmentation (3DC [36], SST [17] and AOT [53]), and salient object detection (iGAN [37] and LGVT [55]). These results are collected from the AVS benchmark [57, 56], which are transferred from the original tasks to the AVS tasks.
As shown in Table 1, our AVSegFormer exceeds these methods by large margins. For instance, on the S4 subset, AVSegFormer-R50 achieves an impressive mIoU of 76.45, which is 1.51 points higher than the best LGVT. Although LGVT has a better Swin-T [33] backbone, our AVSegFormer with ResNet-50 backbone still performs better regarding mIoU. In addition, AVSegFormer-PVTv2 produces an outstanding mIoU of 82.06 and an F-score of 89.9 on this subset, which is 7.12 mIoU and 2.6 F-score higher than LGVT, respectively. On the MS3 subset, AVSegFormer-R50 outperforms the best iGAN with 6.64 mIoU and 8.4 F-score, while AVSegFormer-PVTv2 further raised the bar with an exceptional improvement of 15.47 mIoU and 14.9 F-score. On the AVSS subset, our AVSegFormer-R50 yields 24.93 mIoU and 29.3 F-score, which are slightly lower than the best method AOT. Nevertheless, AVSegFormer-PVTv2 obtains an impressive performance of 36.66 mIoU and 42.0 F-score, surpassing AOT by 11.26 mIoU and 11.0 F-score, respectively.
Comparison with AVSBench baseline [57, 56]. Then, we compare our AVSegFormer with the AVSBench baseline, which is the current state-of-the-art method for audio-visual segmentation. As reported in Table 1, on the S4 subset, AVSegFormer-R50 achieves 3.66 mIoU and 1.1 F-score improvements over AVSBench-R50, while AVSegFormer-PVTv2 surpasses AVSBench-PVTv2 by 3.32 mIoU and 2.0 F-score. On the MS3 subset, AVSegFormer-PVTv2 surpasses AVSBench-PVTv2 with a margin of 4.36 mIoU and 4.8 F-score. On the AVSS subset, AVSegFormer-R50 and AVSegFormer-PVTv2 achieve significant results with an mIoU improvement of 4.75 and 6.89, and a substantial F-score improvement of 4.1 and 6.8. These results demonstrate that AVSegFormer outperforms the AVSBench baseline on all sub-tasks, becoming a new state-of-the-art method for audio-visual segmentation.
4.4. Ablation Study
In this section, we conduct ablation experiments to verify the effectiveness of each key design in the proposed AVSegFormer. Specifically, we adopt PVTv2 [49] as the backbone and conduct extensive experiments on the S4 and MS3 sub-tasks. Other training settings are the same as in Section 4.2.
Number of queries. To analyze the impact of the number of queries on the model’s performance, we conducted experiments with varying numbers of queries for the decoder input, specifically 1, 100, 200, and 300. Our results reveal a positive correlation between the number of queries and the model performance, with the optimal performance obtained when the number of queries was set to 300. Table 2a presents these findings.
Effect of learnable queries. We further investigated the impact of learnable queries in the decoder inputs. As shown in Table 2b, the improvement due to the learnable queries is relatively small in the single sound source task (S4), while it brings significant improvement in the multiple sound source task (MS3). This can be attributed to the complexity of sounding objects. In S4, since there is only one sounding object that remains unchanged, its auditory features are distinct enough. In contrast, in MS3, it becomes relatively difficult for the model to learn target features solely based on audio features, due to the mixing and change of sound sources. The learnable queries partially compensate for this limitation. The results indicate that although the concept of learnable queries is simple, it can produce excellent results in complex scenarios.
Effect of audio-visual mixer. We then studied the impact of the audio-visual mixer in our model. There are two versions designed for this module, as illustrated in Figure 4. The cross-attention mixer (CRA) utilizes visual features as queries and audio features as keys/values for cross-attention, aiming to bring audio information into visual features in the early stages. The channel-attention mixer (CHA) introduced the mechanism of channel attention with audio features as queries and visual features as keys/values. As presented in Table 2c, the design of CHA brought greater performance improvement compared to CRA.
In addition, to validate our suppose that the audio-visual mixer can amplify relevant features while suppressing irrelevant ones, we also visualize features before and after the audio-visual mixer along with their predicted masks. Figure 6 presents the visualization results. We compare the original features with the mixed features generated by the audio-visual mixer and selected 9 channels for rendering. It is evident that for the sounding object (right girl and dog), the mixer effectively enhanced its features. Meanwhile, the non-sounding objects (left girl, table/bed, or background) experienced some degree of suppression. As a result, the predicted mask without mixer may segment out the wrong target (left girl), and the mixer can lead to a more accurate prediction of the correct sounding object. These findings align with our hypothesis and further substantiate the effectiveness of the audio-visual mixer.
Effect of intermediate mask loss. We finally conducted experiments to learn the impact of the intermediate mask loss. Similarly, we designed two versions for the mask reformer module as shown in Figure 5. In the attention-only reformer (AR), we directly feed the attention map of each cross-attention layer into an FPN network to generate intermediate masks. In the query-mixed reformer (QR), we multiply the attention map with the output queries of the corresponding decoder layer, and then feed the generated features into the FPN. The experimental results presented in Table 2d demonstrate that both versions have brought some performance improvement, with the adopted QR outperforming AR.
Qualitative analysis. We also present the visualization results of AVSegFormer compared with those of AVSBench on three audio-visual segmentation tasks in Figure 7. The top line of each task displays the raw images, and the second line displays the ground truth. The last two lines display the predicted masks by AVSBench and AVSegFormer, respectively. The visualization results clearly demonstrate that our method performs better than the previous method. It has a strong ability in target localization and semantic understanding, which can accurately segment and classify the sounding objects in each task. Furthermore, in multiple sound sources scenes, the model can effectively identify the correct sound source and accurately segment the target object. These results highlight the effectiveness and robustness of our method.
5. Conclusion
In this paper, we propose AVSegFormer, a novel framework that leverages the power of transformers to achieve leading performance in audio-visual segmentation tasks. First, our method introduces learnable and audio queries, enabling our network to dynamically attend to relevant visual features and significantly enhance segmentation performance. Second, we design an audio-visual mixer that selectively amplifies or suppresses different visual channels, making the visual features better adapt to diverse audio inputs. Additionally, we propose an intermediate mask loss to improve the effectiveness of the training process. Our experimental results demonstrate the superior performance of AVSegFormer compared to existing state-of-the-art methods, and a series of ablation studies validate the effectiveness of our proposed components.
Appendix A Appendix
We listed the detailed settings of our models in Table 3.
| Settings | S4 | MS3 | AVSS |
| input resolution | |||
| frames | 5 | 5 | 10 |
| embedding dimension | 256 | 256 | 256 |
| transformer layers | 6 | 6 | 6 |
| number of queries | 300 | 300 | 300 |
| mask loss coefficient | 0.1 | 0.1 | 0.1 |
| learnable queries | ✓ | ✓ | ✓ |
| audio-visual mixer | ✓ | ✓ | ✓ |
| batch size | 4 | 4 | 4 |
| optimizer | AdamW | AdamW | AdamW |
| learning rate | |||
| drop path rate | 0.1 | 0.1 | 0.1 |
| epoch | 30 | 60 | 30 |

### v4

@@
 AVSegFormer: Audio-Visual Segmentation with Transformer
-Abstract.
-The combination of audio and vision has long been a topic of interest in the multi-modal community.
-Recently, a new audio-visual segmentation (AVS) task has been introduced, aiming to locate and segment the sounding objects in a given video.
-This task demands audio-driven pixel-level scene understanding for the first time, posing significant challenges.
-In this paper, we propose AVSegFormer, a novel framework for AVS tasks that leverages the transformer architecture.
-Specifically, we introduce audio queries and learnable queries into the transformer decoder, enabling the network to selectively attend to interested visual features.
-Besides, we present an audio-visual mixer, which can dynamically adjust visual features by amplifying relevant and suppressing irrelevant spatial channels.
-Additionally, we devise an intermediate mask loss to enhance the supervision of the decoder, encouraging the network to produce more accurate intermediate predictions.
+Abstract
+Audio-visual segmentation (AVS) aims to locate and segment the sounding objects in a given video, which demands audio-driven pixel-level scene understanding.
+The existing methods cannot fully process the fine-grained correlations between audio and visual cues across various situations dynamically.
+They also face challenges in adapting to complex scenarios, such as evolving audio, the coexistence of multiple objects, and more.
+In this paper, we propose AVSegFormer, a novel framework for AVS that leverages the transformer architecture.
+Specifically, It comprises a dense audio-visual mixer, which can dynamically adjust interested visual features, and a sparse audio-visual decoder, which implicitly separates audio sources and automatically matches optimal visual features.
+Combining both components provides a more robust bidirectional conditional multi-modal representation, improving the segmentation performance in different scenarios.
 Extensive experiments demonstrate that AVSegFormer achieves state-of-the-art results on the AVS benchmark.
 The code is available at https://github.com/vvvb-github/AVSegFormer.
-1. Introduction
-Audio and vision are closely intertwined modalities that play crucial roles in the perception of the world.
-For example, we rely heavily on auditory and visual cues to comprehend and navigate our surroundings.
-These underlying connections motivated many audio-visual tasks, such as audio-visual correspondence [4, 5], audio-visual event localization [29, 31, 46], audio-visual video parsing [45, 51, 30], and sound source localization [4, 5, 43].
-However, due to the lack of pixel-level annotations for these tasks, they are often limited to the frame/temporal level, which makes them become audio-informed image classification problems eventually.
-Recently, a novel audio-visual segmentation (AVS) [56] task has been proposed, which aims to segment sounding objects from video frames corresponding to a given audio.
-It includes three sub-tasks: single sound source segmentation (S4), multiple sound source segmentation (MS3), and audio-visual semantic segmentation (AVSS).
+Introduction
+Just as humans effortlessly establish meaningful connections between audio and visual signals, capturing the rich information they convey, the intertwined modalities of audio and vision play pivotal roles in observing and comprehending the real world.
+Based on this insight, a wide range of audio-visual understanding tasks, such as audio-visual correspondence (Arandjelovic and Zisserman 2017, 2018), audio-visual event localization (Lin, Li, and Wang 2019; Lin and Wang 2020), audio-visual video parsing (Tian, Li, and Xu 2020; Wu and Yang 2021), and sound source localization (Arandjelovic and Zisserman 2017, 2018) have been proposed and actively explored in recent research.
+Unlike these coarse-grained tasks, audio-visual segmentation (AVS) (Zhou et al.
+2023) proposes more fine-grained perceptive goals, aiming to locate the audible frames and delineate the shape of the sounding objects (Zhou et al.
+2022, 2023).
+To be more specific, this task involves three sub-tasks: single sound source segmentation (S4), multiple sound source segmentation (MS3), and audio-visual semantic segmentation (AVSS).
 Figure 1 illustrates the objectives of the three sub-tasks.
-Compared to previous audio-visual tasks, the AVS task presents a greater challenge as it requires the network to not only locate the audible frames but also delineate the shape of the sounding objects [57, 56].
-This demands the alignment of multiple modalities and necessitates a detailed understanding of the scenarios.
-These unique characteristics of the AVS task make many existing methods [7, 39, 36] designed for other audio-visual tasks may be sub-optimal on AVS.
-Therefore, designing new methods tailored for AVS becomes necessary.
-For the brand-new AVS task, AVSBench [56] designed a strong baseline method that achieves state-of-the-art audio-visual segmentation performance.
-Figure 2(a) illustrates its network architecture, which incorporates a modality fusion module before the convolution-based decoder (i.e., Semantic FPN [26]) to enable audio-visual segmentation.
-This design is simple yet effective, but some inherent flaws of convolution still limit it.
-(1) The effective receptive field of convolutions is relatively small.
-Even with a deep decoder, the audio feature still cannot capture long-range visual dependencies, which restricts its performance.
-(2) Convolution is an operator with static weights, which is difficult to provide different visual features conditioned by the input audio.
-To remedy these issues, we propose AVSegFormer, a novel framework for audio-visual segmentation with transformers.
+Their fine-grained perceptual goal demands the model to possess the capability to discern the intricate relationship between each image pixel and audio information.
+However, the existing methods (Chen et al.
+2021; Qian et al.
+2020; Mahadevan et al.
+2020) developed for other audio-visual tasks face challenges when directly applied in this context.
+AVSBench (Zhou et al.
+2023) first proposes the fine-grained perceptive method for AVS tasks that achieves state-of-the-art audio-visual segmentation performance.
+Figure 2(a) illustrates its network architecture, which incorporates a modality fusion module before Semantic FPN (Kirillov et al.
+2019) to enable audio-visual segmentation.
+This method is simple and effective yet falls short of fully mining the fine-grained correlations between audio and visual cues across various situations.
+First, a series of targets sound simultaneously in scenes of multiple sound sources.
+The mixed audio signal with higher information density is difficult to attend to the visual signal adaptively.
+Second, the simple-fusion model may make it difficult to extract audio information bound with the corresponding frame when the sound source objects change over time.
+Additionally, in cases where multiple objects coexist within a single contextual frame (e.g., two guitars or one person and one guitar), the vanilla dense segmentation method struggles to untangle the different sound sources in the audio, making it challenging to achieve precise localization and segmentation in a one-to-one manner.
+To remedy these issues, we propose AVSegFormer, a novel framework for audio-visual segmentation with the transformer architecture.
 The brief architecture is shown in Figure 2(b).
-First, we introduce audio queries along with learnable queries into the segmentation decoder, enabling the network to selectively attend to interested visual features.
-Second, we present an audio-visual mixer, which can dynamically adjust visual features by amplifying relevant and suppressing irrelevant spatial channels, allowing visual features to adapt to diverse audio features.
-Third, an intermediate mask loss is designed to enhance the supervision of the decoder, which encourages the network to produce more accurate intermediate predictions and helps refine the final segmentation outputs.
-We evaluate AVSegFormer on the three sub-tasks of AVS, including S4, MS3, and AVSS, with widely-used backbones ResNet-50 [20] and PVTv2 [49].
-Our experimental results demonstrate that AVSegFormer significantly outperforms existing state-of-the-art methods, such as LGVT [55], SST [17], iGAN [37], and AVSBench baseline [57].
-Specifically, AVSegFormer-R50 achieves 76.45, 49.53, and 24.93 mIoU on S4, MS3, and AVSS, surpassing the AVSBench-R50 by 3.66, 1.65, and 4.75 mIoU, respectively.
-Furthermore, using PVTv2 as the backbone, AVSegFormer yields consistently higher segmentation performance on all three sub-tasks, setting new state-of-the-art records of 82.06, 58.36, and 36.66 mIoU.
-Overall, our contributions to this work are four-fold.
-We employ audio queries and learnable queries to selectively attend to relevant visual features, overcoming the limitations of previous convolutional-based approaches.
-We design an audio-visual mixer, which can amplify relevant visual features and suppress irrelevant visual features in response to audio cues.
-We present an intermediate mask loss that provides additional supervision, encouraging more accurate intermediate results and improving final prediction.
-We conduct extensive experiments on three sub-tasks of AVS. These results demonstrate that our method achieves state-of-the-art performance on the AVS benchmark.
-2. Related Works
-2.1. Multi-Modal Tasks
-In recent years, multi-modal tasks have gained significant attention in the research community due to their potential for advancing the understanding of complex real-world scenarios.
-These tasks aim to exploit the complementary information from different modalities, such as vision, audio, and text, to improve the overall performance of various applications.
+AVSegFormer comprises four key components: (1) a transformer encoder building the mask feature, (2) an audio-visual mixer generating the vision-conditioned mask feature, (3) a query generator initializing sparse audio-conditioned queries, and (4) a cross-modal transformer decoder separating potential sparse object in the visual feature.
+Among them, we design an auxiliary mixing loss to supervise the cross-modal feature generated by the audio-visual mixer.
+It encourages the model to attend to useful information within complex audio semantics and predict the segmentation mask densely.
+Compared to the dense vision-conditioned mixers, the cross-modal transformer decoder aims to build potential sparse queries with the reversed condition.
+It implicitly separates audio sources and automatically matches optimal visual features for different queries.
+These queries will be combined with the vision-conditioned feature map through the matrix production.
+At last, the fused bidirectional conditional feature maps will be used for the final segmentation.
+Overall, our contributions to this work are three-fold:
+(1) We propose AVSegFormer, a transformer-based method for three scenarios of the audio-visual segmentation tasks.
+It combines bidirectional conditional cross-modal feature fusion to provide a more robust audio-visual segmentation representation.
+(2) We propose a dense audio-visual mixer and a sparse audio-visual decoder to provide pixel-level and instance-level complementary representations that can efficiently adapt the scenarios of multiple sound sources and objects.
+(3) Extensive experiments on three sub-tasks of AVS are conducted, demonstrating that AVSegFormer significantly outperforms existing state-of-the-art methods (Mao et al. 2023; Zhou et al. 2022).
+Related Works
+Multi-Modal Tasks
+In recent years, multi-modal tasks have gained significant attention in the research community.
 Among these, text-visual tasks have attracted considerable interest from researchers.
-Numerous works focus on related tasks, such as visual question answering [3, 50, 2, 9] and visual grounding [13, 15, 25, 32].
-In addition to text-visual tasks, audio-visual tasks are emerging as hot spots in the multi-modal field.
-Several related tasks include audio-visual correspondence [4, 5], audio-visual event localization [29, 31, 46], and sound source localization [4, 5, 43].
-Concurrently, some works [59, 48] have proposed unified architectures capable of handling nearly all types of modalities using a consistent format for embeddings.
-Most of these works are based on transformer architecture [47], which demonstrates a strong cross-modal capability.
-Their achievements highlight the reliability of transformers in the multi-modal field.
-As a recently proposed multi-modal task, audio-visual segmentation (AVS) [57, 56] shares many commonalities with the aforementioned tasks.
+Numerous works focus on related tasks, such as visual question answering (Antol et al.
+2015; Wu et al.
+2016) and visual grounding (Deng et al.
+2021; Kamath et al.
+2021).
+In addition to text-visual tasks, audio-visual tasks are emerging as hot spots.
+Related tasks include audio-visual correspondence (Arandjelovic and Zisserman 2017, 2018), audio-visual event localization (Lin, Li, and Wang 2019; Lin and Wang 2020), and sound source localization (Arandjelovic and Zisserman 2017, 2018).
+Concurrently, many works (Zhu et al.
+2022; Wang et al.
+2022a) have proposed unified architectures to deal with multi-modal inputs.
+Most of these works are based on transformer architecture (Vaswani et al.
+2017), demonstrating a strong cross-modal capability.
+Their success highlights the reliability of transformers in the multi-modal field.
+As a recently proposed multi-modal task, audio-visual segmentation (Zhou et al.
+2022, 2023) shares many commonalities with the aforementioned tasks.
 The pioneering works in these areas have significantly inspired our research of AVSegFormer.
-2.2. Sound Source Localization
-Sound source localization (SSL) is an important problem in the audio-visual multi-modal community and is also the most related task to audio-visual segmentation.
-It aims to locate the regions for sound sources in a video sequence, but the results are usually represented as heat maps.
-The major challenge in the SSL problem is dealing with multiple sound sources.
-In prior arts, Hu et al.
-[22] divided audio-visual features into multiple clustering centers and used center distance as a supervised signal to locate paired audio-visual information.
-Qian et al.
-[39] trained an audio-visual correspondence model to extract coarse feature representations of audio and visual signals and used Grad-CAM [42] to locate specific categories of features.
-Hu et al.
-[23] adopted a two-stage approach by first learning audio-visual semantics under a single sound source and then using this knowledge to help locate multiple sound sources.
-Besides, Rouditchenko et al.
-[41] tackled the problem by unraveling the concept of categories in neural networks.
-Although these SSL methods indicate which areas in the image emit sound, they do not clearly depict the shape of the object, which is another challenge in the AVS task.
-Furthermore, the above methods all rely on unsupervised learning when capturing the shape of the detected object, which may result in inaccurate localization.
-Nevertheless, handling multiple sound sources in SSL methods offers valuable insights and references for our work.
-2.3. Vision Transformer
-During the past few years, Transformer [47] has experienced rapid development in natural language processing.
-Following this success, the Vision Transformer (ViT) [16] emerged, bringing the transformer into the realm of computer vision and yielding impressive results.
-Numerous works [33, 49, 10] have built upon ViT, leading to the maturation of vision transformers.
+Vision Transformer
+During the past few years, Transformer (Vaswani et al.
+2017) has experienced rapid development in natural language processing.
+Following this success, the Vision Transformer (ViT) (Dosovitskiy et al.
+2020) emerged, bringing the transformer into the realm of computer vision and yielding impressive results.
+Numerous works (Liu et al.
+2021; Wang et al.
+2022b; Chen et al.
+2022) have built upon ViT, leading to the maturation of vision transformers, especially in object detection and image segmentation tasks.
 As the performance of vision transformers continues to advance, they are increasingly replacing CNNs as the mainstream paradigm in the field of computer vision, especially in object detection and image segmentation tasks.
-Carion et al.
-[6] proposed the DETR model and designed a novel bipartite matching loss based on the transformer architecture, paving the way for new research directions in vision transformers.
-Subsequently, improved frameworks such as Deformable DETR [58] and DINO [54] are proposed, introducing mechanisms like deformable attention and denoise training.
+For downstream tasks, Carion et al.
+(2020) proposed the DETR model and designed a novel bipartite matching loss based on the transformer architecture.
+Subsequently, improved frameworks such as Deformable DETR (Zhu et al.
+2020) and DINO (Zhang et al.
+2022) are proposed, introducing mechanisms like deformable attention and denoise training.
 These arts take vision transformers to new heights.
 The remarkable performance of vision transformers has also inspired us to apply this paradigm to AVS tasks, anticipating further advancements in the field.
-2.4. Image Segmentation
+Image Segmentation
 Image segmentation is a critical visual task that involves partitioning an image into distinct segments or regions.
 It includes three different tasks: instance segmentation, semantic segmentation, and panoptic segmentation.
-Instance segmentation predicts the mask of each object instance and its corresponding category, while semantic segmentation needs to classify each pixel in the image into different semantic categories.
-Panoptic segmentation unifies instance and semantic segmentation tasks and predicts the mask of each object instance or background segment.
-Early research proposed specialized models for these tasks, such as Mask R-CNN [19] and HTC [8] for instance segmentation, or FCN [34] and U-Net [40] for semantic segmentation.
-After panoptic segmentation was proposed, some related research [26, 52, 28] were conducted and designed universal models for both tasks.
+Early research proposed specialized models for these tasks, such as Mask R-CNN (He et al.
+2017) and HTC (Chen et al.
+2019) for instance segmentation, or FCN (Long, Shelhamer, and Darrell 2015) and U-Net (Ronneberger, Fischer, and Brox 2015) for semantic segmentation.
+After panoptic segmentation was proposed, some related research (Kirillov et al.
+2019; Xiong et al.
+2019; Li et al.
+2022b) were conducted and designed universal models for both tasks.
 The recent introduction of the transformer has led to the development of new models that can unify all the segmentation tasks.
-Mask2Former [11] is one such model that introduces mask attention into the transformer and improves MaskFormer [12].
-Mask DINO [27] is a unified transformer-based framework for both detection and segmentation.
-Recently, OneFormer [24] presented a new multi-task universal image segmentation framework with transformers.
-These models have brought image segmentation to a new level, demonstrating the potential of transformer architecture in vision tasks.
+Mask2Former (Cheng et al.
+2022) is one such model that introduces mask attention into the transformer.
+Mask DINO (Li et al.
+2022a) is a unified transformer-based framework for both detection and segmentation.
+Recently, OneFormer (Jain et al.
+2022) presented a new universal image segmentation framework with transformers.
+These models have brought image segmentation to a new level.
 Considering that the AVS task involves segmentation, these methods have significantly contributed to our work.
-3. METHOD
-3.1. Overall Architecture
+Methods
+Overall Architecture
 Figure 3 illustrates the overall architecture of our method.
-In contrast to previous CNN-based segmentation methods [57, 56], we design a query-based framework to leverage the transformer architecture.
-Specifically, our model combines audio queries with learnable queries, allowing it to adjust its focus on visual features dynamically.
-Additionally, we design an audio-visual mixer and an intermediate mask loss as auxiliary components.
-The audio-visual mixer aids in amplifying relevant features and suppressing irrelevant ones, while the intermediate mask loss helps supervise intermediate predictions for enhancing performance.
-The pipeline of AVSegFormer consists of four stages.
-First, a visual backbone and an audio backbone are employed to extract features from video and audio frames, respectively.
-Second, the transformer encoder refines the visual features and generates an initial mask feature, which serves as the basis for predicting the final mask.
-Third, an audio-visual mixer is utilized to amplify feature channels relevant to sounding objects while suppressing those that are irrelevant.
-Lastly, the transformer decoder incorporates audio queries and learnable queries, capturing richer features about sounding objects and predicting the final mask.
-3.2. Multi-Modal Representation
+In contrast to previous CNN-based methods (Zhou et al.
+2022, 2023), we design a query-based framework to leverage the transformer architecture.
+Specifically, the query generator initializes audio queries, and the transformer encoder extracts multi-scale features, which serve as inputs of the transformer decoder for separating potential sparse objects.
+Besides, the audio-visual mixer will further amplify relevant features and suppress irrelevant ones, while the auxiliary mixing loss helps supervise the reinforced features.
+Multi-Modal Representation
 Visual encoder.
-We follow the feature extraction process adopted in previous methods [57, 56], which involves using a visual backbone and an audio backbone to extract video and audio features, respectively.
-For videos, the dataset provides pre-extracted frame images, making the process similar to image feature extraction.
+We follow the feature extraction process adopted in previous methods (Zhou et al.
+2022, 2023), which uses a visual backbone and an audio backbone to extract video and audio features, respectively.
+The dataset provides pre-extracted frame images from videos, making the process similar to image feature extraction.
 Specifically, the input video frames are denoted as , in which denotes the number of frames.
-Then, we use a visual backbone (e.g., ResNet-50 [20]) to extract hierarchical visual features , which can be written as:
+Then, we use a visual backbone (e.g., ResNet-50 (He et al.
+2016)) to extract hierarchical visual features , which can be written as:
 in which and .
-represents the number of output channels of the stage of the visual backbone.
-In other words, is a list of multi-scale features, where each feature map is half resolution of the previous one.
 Audio encoder.
-The process of audio feature extraction follows the VGGish [21] method.
+The process of audio feature extraction follows the VGGish (Hershey et al.
+2017) method.
 Initially, the audio is resampled to 16kHz mono audio , where is related to the audio duration.
 Then, we perform a short-time Fourier transform to obtain a mel spectrum.
-The mel spectrum is calculated by mapping the spectrum to a 64th-order mel filter bank and subsequently fed into the VGGish model to obtain the audio features , where represents the number of frames and defaults to 128 in the VGGish.
-Feature transformation.
-For the convenience of subsequent usage, we use multiple linear layers to unify the number of channels for all features.
-Specifically, all features extracted by the visual backbone and audio backbone are transformed into dimensions, which also equals the embedding dimension of the transformer encoder and decoder.
-Typically, is set to 256 by default.
-3.3. Transformer Encoder
-The transformer encoder is responsible for multi-scale feature fusion and mask feature generation.
-Specifically, we collect the visual features of three resolutions (i.e., 1/8, 1/16, and 1/32), and then flatten and concatenate them as the input for the transformer encoder.
-After that, the output features of the transformer encoder are reshaped to their original shapes, and the 1/8-scale features are taken out separately and 2 upsampled.
-Finally, we add the upsampled features to the 1/4-scale features from the visual backbone and obtain the initial mask features , where , and is the embedding dimension of the transformer encoder and decoder.
-3.4. Audio-Visual Mixer
-As illustrated in Figure 3, the segmentation mask is generated based on the mask feature, which plays a crucial role in the final prediction results.
-However, since the audio features can vary widely, a static network may not be able to capture all of the relevant information.
-This limitation may hinder the model’s ability to identify potential sounding objects accurately.
+The mel spectrum is calculated by mapping the spectrum to a 64th-order mel filter bank and then fed into the VGGish model to obtain the audio features , where means the number of frames.
+Query Generator
+The query generator is designed to generate sparse audio-conditioned queries, which helps the model fully understand the auditory information.
+At the beginning, we have an initial query and audio feature , here represents the number of queries.
+We employ as queries and as keys and values, feed them into the query generator, and obtain the audio query .
+Finally, we incorporate audio query and learnable query as a mixed query for the input of the transformer decoder.
+The addition of learnable queries enhances our model’s adaptability for various AVS tasks and datasets.
+It enables the model to learn dataset-level contextual information, and adjust the attention allocated to different sounding targets.
+Transformer Encoder
+The transformer encoder is responsible for building the mask feature.
+Specifically, we collect the backbone features of three resolutions (i.e., 1/8, 1/16, and 1/32), and then flatten and concatenate them as the input queries for the transformer encoder.
+After that, the output features are reshaped to their original shapes, and the 1/8-scale features are taken out separately and 2 upsampled.
+Finally, we add the upsampled features to the 1/4-scale features from the visual backbone and obtain the mask feature , where , and is the embed dimension.
+Audio-Visual Mixer
+As illustrated in Figure 3, the segmentation mask is generated based on the mask feature, which plays a crucial role in the final results.
+However, since the audio semantics can vary widely, a static network may not be able to capture all of the relevant information.
+This limitation may hinder the model’s ability to identify inconspicuous sounding objects.
 To address this issue, we propose an audio-visual mixer as shown in Figure 4(b).
-The design of this module is based on channel attention, which allows the model to selectively amplify or suppress different visual channels depending on the audio input, improving its ability to capture complex audio-visual relationships.
+The design of this module is based on channel attention, which allows the model to selectively amplify or suppress different visual channels depending on the audio feature, improving its ability to capture complex audio-visual relationships.
 Specifically, the mixer learns a set of weights through audio-visual cross-attention, and applies them to highlight the relevant channels.
 The whole process can be represented as follows:
-Here, and represent the input audio features and the initial mask features, and denotes the mixed mask features.
-Besides, means the number of attention heads, which is set to 8 by default following common practice.
-3.5. Transformer Decoder
-Audio query.
-The transformer decoder is designed to learn semantic-rich features of the sounding objects.
-We repeat the audio feature to the number of queries , and employ it as the audio queries .
-As the decoding process continues, the queries continuously aggregate visual information from the encoder’s outputs.
-The output queries ultimately combine the auditory and visual modalities and contain richer features of the sounding objects.
-Learnable query.
-To enhance the model’s ability to capture in-depth information about the sounding object, we add learnable queries to the audio queries .
-The learnable queries enhance our model’s adaptability for various AVS tasks and datasets.
-Specifically, it enables the model to learn dataset-level contextual information, and adjust the attention allocated to different target categories.
-Furthermore, learnable queries empower the model with a more robust capability to process target semantics, ultimately improving segmentation accuracy.
-Mask generation.
-To generate the segmentation masks, we multiply the mask feature obtained from the audio-visual mixer with the output queries from the decoder.
+Here, and represent the input audio feature and the initial mask feature, and denotes the mixed mask feature.
+means the number of attention heads, which is set to 8 by default following common practice.
+Transformer Decoder
+The transformer decoder is designed to build potential sparse queries, and optimally match the visual features with corresponding queries.
+We utilize the mixed query as the input query and the multi-scale visual features as key/value.
+As the decoding process continues, the output queries continuously aggregate with visual features, ultimately combining the auditory and visual modalities and containing various target information.
+To generate the segmentation masks, we multiply the mask feature obtained from the audio-visual mixer with the mixed queries from the query generator.
 Then, an MLP is used to integrate different channels.
-Additionally, we introduce a residual connection to ensure that the fusion of auditory information does not result in excessive loss of visual information.
-Finally, the model predicts the segmentation mask through a simple linear layer:
-Here, represents the MLP process, and means the linear layer. The output is the predicted segmentation mask, with the dimension denotes the number of semantic classes.
-3.6. Loss Function
-Intermediate mask loss.
-With the introduction of the deep transformer decoder, achieving satisfying performance only by supervising the final prediction becomes more challenging.
-To address this issue, we design an intermediate mask loss as the auxiliary loss, which supervises each layer of the transformer decoder.
-Specifically, the intermediate mask loss is based on the cross-attention operation of each decoder layer.
-First, we utilize the input queries and keys to compute the attention map .
-Then, we multiply the attention map with the output queries of the current decoder layer and feed it through a simple FPN [44] to predict a mask , as shown in Figure 5(c).
-Finally, we calculate the Dice loss [38] between this predicted mask and the ground truth to obtain the intermediate mask loss .
+Finally, the model predicts the mask through a fully connected layer:
+Here, represents the MLP layer, and means the fully connected layer. The output is the predicted segmentation mask, with the dimension denotes the number of semantic classes.
+Loss Function
+Auxiliary mixing loss.
+With the introduction of the audio-visual mixer, our model has maintained great capability in dealing with plentiful audio semantics, but its robustness is still insufficient when facing complex scenes.
+Thus, we design a mixing loss to supervise the mixer, enabling the model to more accurately locate target objects.
+Specifically, we integrate all channels of the mask feature through a linear layer and predict a binary mask.
+At the same time, we extract all foreground labels in the ground truth as a new binary label and calculate the Dice loss (Milletari, Navab, and Ahmadi 2016) between them.
 Total loss.
-The total loss function comprises two parts: IoU loss and mask loss.
-As shown in Figure 5(a), the IoU loss is calculated by comparing the final segmentation mask with the ground truth.
-Here, we also use Dice loss [38] for supervision.
+The loss function comprises two parts: IoU loss and mixing loss.
+The IoU loss is calculated by comparing the final segmentation mask with the ground truth.
+Here, we use Dice loss (Milletari, Navab, and Ahmadi 2016) for supervision.
 Considering that in AVS tasks, the proportion of segmented objects occupying the entire image is relatively small, the model can better focus on the foreground and reduce interference from the background by using Dice loss.
-Thus, the total loss of our method is as follows:
-Here, represents the number of decoder layers, and is a coefficient that controls the effect of the auxiliary loss. We choose in our experiments as it performs best.
+Thus, the total loss of our method is:
+Here, is a coefficient that controls the effect of the auxiliary loss. We set as it performs best.
+Discussion
+AVSegFormer adopts a framework for segmentation that resembles Mask2Former (Cheng et al.
+2022).
+However, it distinguishes itself by tailoring enhancements specifically for AVS tasks, accommodating the input of multi-modal information, which is a capability not inherent in Mask2Former.
+We introduce dual-tower backbone networks to ensure comprehensive extraction of both visual and auditory features.
+Besides, we devise a novel dense audio-visual mixer and a sparse audio-visual decoder that empower the proposed AVSegFormer to leverage auditory cues effectively, resulting in enhanced segmentation performance.
+These tailored designs hold substantial implications for addressing the emerging challenges of the AVS tasks.
+Experiments
 Method | Backbone | S4 | MS3 | AVSS | Reference | |||
 F-score | mIoU | F-score | mIoU | F-score | mIoU | |||
-| LVS [7] | ResNet-50 [20] | 51.0 | 37.94 | 33.0 | 29.45 | CVPR’2021 | ||
-| MSSL [39] | ResNet-18 [20] | 66.3 | 44.89 | 36.3 | 26.13 | ECCV’2020 | ||
-| 3DC [36] | ResNet-34 [20] | 75.9 | 57.10 | 50.3 | 36.92 | 21.6 | 17.27 | BMVC’2020 |
-| SST [17] | ResNet-101 [20] | 80.1 | 66.29 | 57.2 | 42.57 | CVPR’2021 | ||
-| AOT [53] | Swin-B [33] | 31.0 | 25.40 | NeurIPS’2021 | ||||
-| iGAN [37] | Swin-T [33] | 77.8 | 61.59 | 54.4 | 42.89 | ArXiv’2022 | ||
-| LGVT [55] | Swin-T [33] | 87.3 | 74.94 | 59.3 | 40.71 | NeurIPS’2021 | ||
-| AVSBench-R50 [56] | ResNet-50 [20] | 84.8 | 72.79 | 57.8 | 47.88 | 25.2 | 20.18 | ECCV’2022 |
-| AVSegFormer-R50 (ours) | ResNet-50 [20] | 85.9 | 76.45 | 62.8 | 49.53 | 29.3 | 24.93 | |
-| AVSBench-PVTv2 [56] | PVTv2 [49] | 87.9 | 78.74 | 64.5 | 54.00 | 35.2 | 29.77 | ECCV’2022 |
-| AVSegFormer-PVTv2 (ours) | PVTv2 [49] | 89.9 | 82.06 | 69.3 | 58.36 | 42.0 | 36.66 | |
+LVS | ResNet-50 | 51.0 | 37.94 | 33.0 | 29.45 | CVPR’2021 | ||
+MSSL | ResNet-18 | 66.3 | 44.89 | 36.3 | 26.13 | ECCV’2020 | ||
+3DC | ResNet-34 | 75.9 | 57.10 | 50.3 | 36.92 | 21.6 | 17.27 | BMVC’2020 |
+SST | ResNet-101 | 80.1 | 66.29 | 57.2 | 42.57 | CVPR’2021 | ||
+AOT | Swin-B | 31.0 | 25.40 | NeurIPS’2021 | ||||
+iGAN | Swin-T | 77.8 | 61.59 | 54.4 | 42.89 | ArXiv’2022 | ||
+LGVT | Swin-T | 87.3 | 74.94 | 59.3 | 40.71 | NeurIPS’2021 | ||
+AVSBench-R50 | ResNet-50 | 84.8 | 72.79 | 57.8 | 47.88 | 25.2 | 20.18 | ECCV’2022 |
+DiffusionAVS-R50 | ResNet-50 | 86.9 | 75.80 | 62.1 | 49.77 | ArXiv’2023 | ||
+AVSegFormer-R50 (ours) | ResNet-50 | 85.9 | 76.45 | 62.8 | 49.53 | 29.3 | 24.93 | AAAI’2024 |
+AVSegFormer-R50* (ours) | ResNet-50 | 86.7 | 76.38 | 65.6 | 53.81 | 31.5 | 26.58 | AAAI’2024 |
+AVSBench-PVTv2 | PVTv2 | 87.9 | 78.74 | 64.5 | 54.00 | 35.2 | 29.77 | ECCV’2022 |
+DiffusionAVS-PVTv2 | PVTv2 | 90.2 | 81.38 | 70.9 | 58.18 | ArXiv’2023 | ||
+AVSegFormer-PVTv2 (ours) | PVTv2 | 89.9 | 82.06 | 69.3 | 58.36 | 42.0 | 36.66 | AAAI’2024 |
+AVSegFormer-PVTv2* (ours) | PVTv2 | 90.5 | 83.06 | 73.0 | 61.33 | 42.8 | 37.31 | AAAI’2024 |
 S4 | MS3 | |||
 mIoU | F | mIoU | F | |
-| 1 | 81.1 | 88.6 | 57.1 | 67.7 |
-| 100 | 81.4 | 88.9 | 57.5 | 68.2 |
-| 200 | 81.9 | 89.6 | 58.0 | 68.4 |
-| 300 | 82.0 | 89.9 | 58.3 | 69.3 |
+1 | 79.6 | 86.6 | 59.1 | 69.7 |
+100 | 81.4 | 88.9 | 60.5 | 71.2 |
+200 | 82.3 | 89.6 | 61.0 | 72.4 |
+300 | 83.1 | 90.5 | 61.3 | 73.0 |
 learnable | S4 | MS3 | ||
 queries | mIoU | F | mIoU | F |
-| ✓ | 82.0 | 89.9 | 58.3 | 69.3 |
-| 81.8 | 89.6 | 57.2 | 67.9 | |
+✓ | 83.1 | 90.5 | 61.3 | 73.0 |
+82.7 | 89.9 | 58.5 | 70.9 |
 mixer | S4 | MS3 | ||
 mIoU | F | mIoU | F | |
-| 81.4 | 89.0 | 57.4 | 67.9 | |
-| CRA | 81.7 | 89.7 | 57.8 | 68.2 |
-| CHA | 82.0 | 89.9 | 58.3 | 69.3 |
-| mask | S4 | MS3 | ||
-| reformer | mIoU | F | mIoU | F |
-| 81.1 | 88.4 | 57.3 | 68.3 | |
-| AR | 81.7 | 89.6 | 57.9 | 68.8 |
-| QR | 82.0 | 89.9 | 58.3 | 69.3 |
-4. Experiments
-4.1. Dataset
-AVSBench-Object [57] is an audio-visual dataset specifically designed for the audio-visual segmentation task, containing pixel-level annotations.
+81.4 | 88.0 | 59.4 | 70.9 | |
+CRA | 82.7 | 89.7 | 59.8 | 72.2 |
+CHA | 83.1 | 90.5 | 61.3 | 73.0 |
+mix | S4 | MS3 | ||
+loss | mIoU | F | mIoU | F |
+✓ | 83.1 | 90.5 | 61.3 | 73.0 |
+81.3 | 89.1 | 59.6 | 71.3 |
+Dataset
+AVSBench-Object (Zhou et al.
+2022) is an audio-visual dataset specifically designed for the audio-visual segmentation task, containing pixel-level annotations.
 The videos are downloaded from YouTube and cropped to 5 seconds, with one frame per second extracted for segmentation.
-The dataset includes two subsets: a semi-supervised single sound source subset for single sound source segmentation (S4), and a fully supervised multi-source subset for multiple sound source segmentation (MS3).
-S4 subset: The S4 subset contains a total of videos, with videos for training, for validation, and for testing.
-The target objects cover 23 different categories, including humans, animals, vehicles, and musical instruments.
-Besides, this subset is trained in a semi-supervised manner, where each video contains five frames, but only the first frame is annotated.
+The dataset includes two subsets: a single sound source subset for single sound source segmentation (S4), and a multi-source subset for multiple sound source segmentation (MS3).
+S4 subset: The S4 subset contains videos, with videos for training, for validation, and for testing.
+The target objects cover 23 categories, including humans, animals, vehicles, and musical instruments.
 MS3 subset: The MS3 subset includes 424 videos, with 286 training, 64 validation, and 64 testing videos, covering the same categories as the S4 subset.
-Additionally, unlike S4, the MS3 subset is full-supervised with all five frames annotated in training.
-AVSBench-Semantic [56] is an extension of the AVSBench-Object dataset, which offers additional semantic labels that are not available in the original AVSBench-Object dataset.
+AVSBench-Semantic (Zhou et al.
+2023) is an extension of the AVSBench-Object, which offers additional semantic labels that are not available in the original AVSBench-Object dataset.
 It is designed for audio-visual semantic segmentation (AVSS).
 In addition, the videos in AVSBench-Semantic are longer, with a duration of 10 seconds, and 10 frames are extracted from each video for prediction.
-It combines semi-supervised and fully-supervised manners, with labels provided for the first and first five frames of videos inherited from the S4 and MS3 subsets, respectively.
-For the newly collected 7,000 videos, a complete 10-frame label is provided.
 Overall, the AVSBench-Semantic dataset has increased in size by approximately three times compared to the original AVSBench-Object dataset, with 8,498 training, 1,304 validation, and 1,554 test videos.
-Metric. These benchmark datasets employ mean intersection over union (mIoU) and F-score as the evaluation metrics.
-4.2. Implementation Details
-Our method is evaluated on three AVS sub-tasks, including single sound source segmentation (S4), multiple sound source segmentation (MS3), and audio-visual semantic segmentation (AVSS).
-Details.
-We train our AVSegFormer models for the three sub-tasks using an NVIDIA V100 GPU.
-Consistent with previous works [57, 56], we employ AdamW [35] as the optimizer, with a batch size of 4 and an initial learning rate of .
-All video frames are resized to 224 224 resolution.
-For the S4 and MS3 subsets, each video contains 5 frames, while each video in AVSS contains 10 frames.
+Implementation Details
+We train our AVSegFormer models for the three AVS sub-tasks using an NVIDIA V100 GPU.
+Consistent with previous works (Zhou et al.
+2022, 2023), we employ AdamW (Loshchilov and Hutter 2017) as the optimizer, with a batch size of 2 and an initial learning rate of .
 Since the MS3 subset is quite small, we train it for 60 epochs, while the S4 and AVSS subsets are trained for 30 epochs.
-The encoder and decoder in our AVSegFormer both are comprised of 6 layers with an embedding dimension of 256.
-We set the coefficient of the proposed intermediate mask loss to 0.1 for the best performance.
-More detailed training settings can be found in Table 3.
-4.3. Comparison with Prior Arts
-To verify the effectiveness of our method, we conducted a comprehensive comparison between our AVSegFormer and existing methods on the AVS benchmark [57, 56].
-For fairness, we employ the ImageNet-1K [14] pre-trained ResNet-50 [20] or PVTv2 [49] as the backbone to extract visual features, and the AudioSet [18] pre-trained VGGish [21] to extract audio features.
+The encoder and decoder in our AVSegFormer comprise 6 layers with an embedding dimension of 256.
+We set the coefficient of the proposed mixing loss to 0.1 for the best performance.
+More detailed training settings can be found in the supplementary materials.
+Comparison with Prior Arts
+We conducted a comprehensive comparison between our AVSegFormer and existing methods (Zhou et al.
+2022, 2023; Mao et al.
+2023) on the AVS benchmark.
+For fairness, we employ the ImageNet-1K (Deng et al.
+2009) pre-trained ResNet-50 (He et al.
+2016) or PVTv2 (Wang et al.
+2022b) as the backbone to extract visual features, and the AudioSet (Gemmeke et al.
+2017) pre-trained VGGish (Hershey et al.
+2017) to extract audio features.
 Comparison with methods from related tasks.
-Firstly, we compare our AVSegFormer with state-of-the-art methods from three AVS-related tasks, including sound source localization (LVS [7] and MSSL [39]), video object segmentation (3DC [36], SST [17] and AOT [53]), and salient object detection (iGAN [37] and LGVT [55]).
-These results are collected from the AVS benchmark [57, 56], which are transferred from the original tasks to the AVS tasks.
+Firstly, we compare our AVSegFormer with state-of-the-art methods from three AVS-related tasks, including sound source localization (LVS (Chen et al.
+2021) and MSSL (Qian et al.
+2020)), video object segmentation (3DC (Mahadevan et al.
+2020), SST (Duke et al.
+2021) and AOT (Yang, Wei, and Yang 2021)), and salient object detection (iGAN (Mao et al.
+2021) and LGVT (Zhang et al.
+2021)).
+These results are collected from the AVS benchmark (Zhou et al.
+2022), which are transferred from the original tasks to the AVS tasks.
 As shown in Table 1, our AVSegFormer exceeds these methods by large margins.
 For instance, on the S4 subset, AVSegFormer-R50 achieves an impressive mIoU of 76.45, which is 1.51 points higher than the best LGVT.
-Although LGVT has a better Swin-T [33] backbone, our AVSegFormer with ResNet-50 backbone still performs better regarding mIoU.
+Although LGVT has a better Swin-T (Liu et al.
+2021) backbone, our AVSegFormer with ResNet-50 backbone still performs better regarding mIoU.
 In addition, AVSegFormer-PVTv2 produces an outstanding mIoU of 82.06 and an F-score of 89.9 on this subset, which is 7.12 mIoU and 2.6 F-score higher than LGVT, respectively.
 On the MS3 subset, AVSegFormer-R50 outperforms the best iGAN with 6.64 mIoU and 8.4 F-score, while AVSegFormer-PVTv2 further raised the bar with an exceptional improvement of 15.47 mIoU and 14.9 F-score.
-On the AVSS subset, our AVSegFormer-R50 yields 24.93 mIoU and 29.3 F-score, which are slightly lower than the best method AOT.
-Nevertheless, AVSegFormer-PVTv2 obtains an impressive performance of 36.66 mIoU and 42.0 F-score, surpassing AOT by 11.26 mIoU and 11.0 F-score, respectively.
-Comparison with AVSBench baseline [57, 56].
+On the AVSS subset, our AVSegFormer-R50 yields 24.93 mIoU and 29.3 F-score, and AVSegFormer-PVTv2 obtains an impressive performance of 36.66 mIoU and 42.0 F-score, surpassing AOT by 11.26 mIoU and 11.0 F-score, respectively.
+Comparison with AVSBench.
 Then, we compare our AVSegFormer with the AVSBench baseline, which is the current state-of-the-art method for audio-visual segmentation.
 As reported in Table 1, on the S4 subset, AVSegFormer-R50 achieves 3.66 mIoU and 1.1 F-score improvements over AVSBench-R50, while AVSegFormer-PVTv2 surpasses AVSBench-PVTv2 by 3.32 mIoU and 2.0 F-score.
-On the MS3 subset, AVSegFormer-PVTv2 surpasses AVSBench-PVTv2 with a margin of 4.36 mIoU and 4.8 F-score.
+On the MS3 subset, AVSegFormer-PVTv2 surpasses AVSBench-PVTv2 with a margin of 1.65 mIoU and 5.0 F-score.
 On the AVSS subset, AVSegFormer-R50 and AVSegFormer-PVTv2 achieve significant results with an mIoU improvement of 4.75 and 6.89, and a substantial F-score improvement of 4.1 and 6.8.
 These results demonstrate that AVSegFormer outperforms the AVSBench baseline on all sub-tasks, becoming a new state-of-the-art method for audio-visual segmentation.
-4.4. Ablation Study
+Ablation Study
 In this section, we conduct ablation experiments to verify the effectiveness of each key design in the proposed AVSegFormer.
-Specifically, we adopt PVTv2 [49] as the backbone and conduct extensive experiments on the S4 and MS3 sub-tasks.
-Other training settings are the same as in Section 4.2.
+Specifically, we adopt PVTv2 (Wang et al.
+2022b) as the backbone and conduct extensive experiments on the S4 and MS3 sub-tasks.
 Number of queries.
@@
 This can be attributed to the complexity of sounding objects.
-In S4, since there is only one sounding object that remains unchanged, its auditory features are distinct enough.
-In contrast, in MS3, it becomes relatively difficult for the model to learn target features solely based on audio features, due to the mixing and change of sound sources.
-The learnable queries partially compensate for this limitation.
-The results indicate that although the concept of learnable queries is simple, it can produce excellent results in complex scenarios.
+We involve a more detailed discussion in the supplementary materials.
 Effect of audio-visual mixer.
-We then studied the impact of the audio-visual mixer in our model.
-There are two versions designed for this module, as illustrated in Figure 4.
-The cross-attention mixer (CRA) utilizes visual features as queries and audio features as keys/values for cross-attention, aiming to bring audio information into visual features in the early stages.
-The channel-attention mixer (CHA) introduced the mechanism of channel attention with audio features as queries and visual features as keys/values.
+We then studied the impact of the audio-visual mixer on our model.
+Two versions are designed for this module, as illustrated in Figure 4.
+The cross-attention mixer (CRA) utilizes visual features as queries and audio features as keys/values for cross-attention, and the channel-attention mixer (CHA) introduced the mechanism of channel attention with audio features as queries and visual features as keys/values.
 As presented in Table 2c, the design of CHA brought greater performance improvement compared to CRA.
-In addition, to validate our suppose that the audio-visual mixer can amplify relevant features while suppressing irrelevant ones, we also visualize features before and after the audio-visual mixer along with their predicted masks.
-Figure 6 presents the visualization results.
-We compare the original features with the mixed features generated by the audio-visual mixer and selected 9 channels for rendering.
+In addition, we also visualize the mask feature before and after the audio-visual mixer, as shown in Figure 5.
 It is evident that for the sounding object (right girl and dog), the mixer effectively enhanced its features.
 Meanwhile, the non-sounding objects (left girl, table/bed, or background) experienced some degree of suppression.
-As a result, the predicted mask without mixer may segment out the wrong target (left girl), and the mixer can lead to a more accurate prediction of the correct sounding object.
 These findings align with our hypothesis and further substantiate the effectiveness of the audio-visual mixer.
-Effect of intermediate mask loss.
-We finally conducted experiments to learn the impact of the intermediate mask loss.
-Similarly, we designed two versions for the mask reformer module as shown in Figure 5.
-In the attention-only reformer (AR), we directly feed the attention map of each cross-attention layer into an FPN network to generate intermediate masks.
-In the query-mixed reformer (QR), we multiply the attention map with the output queries of the corresponding decoder layer, and then feed the generated features into the FPN.
-The experimental results presented in Table 2d demonstrate that both versions have brought some performance improvement, with the adopted QR outperforming AR.
+Effect of auxiliary mixing loss.
+We finally conducted experiments to learn the impact of the auxiliary mixing loss.
+We train our model with and without the mixing loss, respectively, and report the testing results in Table 2d.
+It is demonstrated that the auxiliary mixing loss can help a lot in the final prediction.
 Qualitative analysis.
 We also present the visualization results of AVSegFormer compared with those of AVSBench on three audio-visual segmentation tasks in Figure 6.
-The top line of each task displays the raw images, and the second line displays the ground truth.
-The last two lines display the predicted masks by AVSBench and AVSegFormer, respectively.
-The visualization results clearly demonstrate that our method performs better than the previous method.
-It has a strong ability in target localization and semantic understanding, which can accurately segment and classify the sounding objects in each task.
-Furthermore, in multiple sound sources scenes, the model can effectively identify the correct sound source and accurately segment the target object.
+The visualization results clearly demonstrate that our method performs better.
+It has a strong ability in target localization and semantic understanding and can effectively identify the correct sound source and accurately segment the target object in multiple sound source scenes.
 These results highlight the effectiveness and robustness of our method.
-5. Conclusion
-In this paper, we propose AVSegFormer, a novel framework that leverages the power of transformers to achieve leading performance in audio-visual segmentation tasks.
-First, our method introduces learnable and audio queries, enabling our network to dynamically attend to relevant visual features and significantly enhance segmentation performance.
-Second, we design an audio-visual mixer that selectively amplifies or suppresses different visual channels, making the visual features better adapt to diverse audio inputs.
-Additionally, we propose an intermediate mask loss to improve the effectiveness of the training process.
-Our experimental results demonstrate the superior performance of AVSegFormer compared to existing state-of-the-art methods, and a series of ablation studies validate the effectiveness of our proposed components.
-Appendix A Appendix
-We listed the detailed settings of our models in Table 3.
-| Settings | S4 | MS3 | AVSS |
-| input resolution | |||
-| frames | 5 | 5 | 10 |
-| embedding dimension | 256 | 256 | 256 |
-| transformer layers | 6 | 6 | 6 |
-| number of queries | 300 | 300 | 300 |
-| mask loss coefficient | 0.1 | 0.1 | 0.1 |
-| learnable queries | ✓ | ✓ | ✓ |
-| audio-visual mixer | ✓ | ✓ | ✓ |
-| batch size | 4 | 4 | 4 |
-| optimizer | AdamW | AdamW | AdamW |
-| learning rate | |||
-| drop path rate | 0.1 | 0.1 | 0.1 |
-| epoch | 30 | 60 | 30 |
+Conclusion
+In this paper, we propose AVSegFormer, a novel audio-visual segmentation framework that leverages the power of transformer architecture to achieve leading performance.
+Specifically, our method comprises four key components: (1) a transformer encoder building the mask feature, (2) a query generator providing audio-conditioned queries, (3) a dense audio-visual mixer dynamically adjusting interested visual features, and (4) a sparse audio-visual decoder separating audio sources and matching optimal visual features.
+These components provide a more robust audio-visual cross-modal representation, improving the AVS performance in different scenarios.
+Extensive experimental results demonstrate the superior performance of AVSegFormer compared to existing state-of-the-art methods.
+Acknowledgement
+This work is supported by the Natural Science Foundation of China under Grant 62372223.